# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.4. Proof of Proposition 9 [Pruned possible graphs for top-level $H^{-1}$]
Proof of Proposition 9.
The proposition follows immediately from Lemmas 51, 52, 53, and 54. ∎
C.5. Proof of Proposition 10 [Extending non-$3$-SIP maps for minimal pairs]
Lemma 55 (Degree $3$ decorations preserve non-$3$-SIP).
Let $G$ be $K_{4}$ and let $H$ be one of its $K_{3}$ subgraphs. Also, let $\ell_{1}$ and $\ell_{2}$ be any squared edge-length maps for $H$ such that some edge is assigned distinct positive real numbers by these maps. If the linkages $(H,\ell_{1})$ and $(H,\ell_{2})$ both have $2$-realizations, then there exist squared edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for $G$ such that
  1. (i)
  2. (ii)
    $\ell^{\prime}_{i}$ agrees with $\ell_{i}$ on the edges of $H$,
  3. (iii)
Proof.
We can clearly choose $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ to satisfy Condition (ii). We show that we can choose these maps to additionally satisfy Conditions (i) and (iii). Let $\{v_{i}\}$ be the vertex set of $G$, let $H$ not contain $v_{4}$, and let $\ell_{1}(v_{1}v_{2})$ and $\ell_{1}(v_{1}v_{2})$ be distinct positive real numbers. Note that there exists a $2$-realization $p$ of $(H,\ell_{1})$ that maps $v_{1}$ to $a_{1}=(0,0)$, $v_{2}$ to $b_{1}=(\ell_{1}(v_{1}v_{2}),0)$, and $v_{3}$ to some point $c_{1}=(x,y)$ with $y\geq 0$. Similarly, there exists a $2$-realization $q$ of $(H,\ell_{2})$ that maps $v_{1}$ to $a_{1}=(0,0)$, $v_{2}$ to $b_{2}=(\ell_{2}(v_{1}v_{2}),0)$, and $v_{3}$ to some point $c_{2}=(w,z)$ with $z\geq 0$. Since $\ell_{1}(v_{1}v_{2})$ and $\ell_{1}(v_{1}v_{2})$ are distinct, there is a unique line $L(b_{1}b_{2})$ that contains $b_{1}$ and $b_{2}$. If $c_{1}=c_{2}$, then consider any point $d$ on the line $\bar{L}(b_{1}b_{2})$ that bisects the line-segment between $b_{1}$ and $b_{2}$. Set $\ell^{\prime}_{i}(v_{1}v_{4})=\|a_{i}-d\|^{2}$, $\ell^{\prime}_{i}(v_{2}v_{4})=\|b_{i}-d\|^{2}$, and $\ell^{\prime}_{i}(v_{3}v_{4})=\|c_{i}-d\|^{2}$. Clearly, $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ satisfy Conditions (i)-(iii), and so the lemma is proved.
Next, if $c_{1}\neq c_{2}$, then there is a unique line $L(c_{1}c_{2})$ that contains $c_{1}$ and $c_{2}$. If $\bar{L}(b_{1}b_{2})$ does not intersect the line $\bar{L}(c_{1}c_{2})$ that bisects the line-segment between $c_{1}$ and $c_{2}$, then a simple algebraic geometry argument shows that we can rotate $b_{1}$ and $c_{1}$ about $(0,0)$ by some angle $\theta$ to obtain the points $b^{\prime}_{1}$ and $c^{\prime}_{1}$, respectively, such that either $c^{\prime}_{1}=c_{2}$ or the line that bisects the line-segment between $b^{\prime}_{1}$ and $b_{2}$ intersects the line that bisects the line-segment between $c^{\prime}_{1}$ and $c_{2}$. In the former case, an argument similar to the one above applies. In the latter case, we can set $d$ to be the intersection of these two lines and then apply a similar argument. This completes the proof. ∎
Proof of Proposition 10.
If $G^{\prime}$ is one of the graphs in Figures 15(a) - 15(c), then it is easy to see that Lemma 24 applies and proves the proposition. Else, if $G^{\prime}$ is the graph in Figure 15(d), then we can first apply Lemma 24 to the subgraph of $G^{\prime}$ obtained by deleting the unlabeled vertices, and then add these vertices back and apply Lemma 55 to prove the proposition. Lastly, if $G^{\prime}$ is the graph in Figure 15(e), then consider the subgraph of $G^{\prime}$ obtained by deleting all but one unlabeled vertex, which we call $x$. We first apply Lemma 24 to this subgraph while treating $xw_{2}$ as $f$, then we apply this lemma again while treating $xw_{2}$ as $f^{\prime}$ and $uw_{1}$ as $f$. Finally, we add the deleted vertices and apply Lemma 55 to prove the proposition. ∎
Appendix D Proof of Theorem 5 [Byproduct: $3$-connected partial $3$-tree star theorem]
The proof of Theorem 5 requires Lemma 20, from Section B.5, and Lemmas 56 and 57, below, which may themselves be independently interesting results about partial $3$-trees.
Lemma 56 (If $G\cup\{wx,wy\}$ has a forbidden minor, then so does $G\cup wx$).
Let $G$ be a $3$-connected graph and $F=\{wx,wy\}$ be some subset of its nonedges. If there exists a forbidden minor $[G\cup F]$ in which $[w]$, $[x]$, and $[y]$ are distinct vertices and $wx$ and $wy$ are retained, then some $f\in F$ is such that $G\cup f$ has a forbidden minor.
Proof.
Let $G^{\prime}=G\cup F$ and assume that the forbidden minor $[G^{\prime}]$ exists. Since $G$ is $3$-connected, it contains at least three internally vertex-disjoint paths $P_{a}$, $P_{b}$, and $P_{c}$ between $w$ and $x$, where $a$, $b$, and $c$ are the first vertices along these paths not contained in $[w]^{-1}$, respectively. Note that $[x]^{-1}\cup[y]^{-1}$ does not contain $a$, $b$, or $c$. Also, since $[w]$ has degree two in $[G^{\prime}]\setminus[F]$, we can choose these paths such that $[a]=[b]$. For any minor $[G^{\prime}]_{\star}$ in which $[w]_{\star}$, $[x]_{\star}$, and $[y]_{\star}$ are distinct vertices and $wx$ and $wy$ are retained, define $a([G^{\prime}]_{\star})$ to be the length of the longest path in $G^{\prime}$ that contains $a$ and consists only of vertices in $V(P_{a})\cap[a]_{\star}^{-1}$. Define $b([G^{\prime}]_{\star})$ similarly. We will show how to obtain a minor $[G^{\prime}]_{A}$ from $[G^{\prime}]$ via a sequence of contractions and component exchanges such that (i) $[G^{\prime}]_{A}$ is a forbidden minor and $[w]_{A}$, $[x]_{A}$, and $[y]_{A}$ are distinct vertices and one of the following statements is true: (ii) either $wx$ or $wy$ is doubled in $[G^{\prime}]_{A}$ or (iii) $a([G^{\prime}]_{A})\geq a([G^{\prime}])$, $b([G^{\prime}]_{A})\geq b([G^{\prime}])$, and at least one of these inequalities is strict. If Statements (i) and (ii) are true, then the lemma is clearly proved. Otherwise, if Statements (i) and (iii) are true, then, since $P_{a}$ and $P_{b}$ are each finite and end with $x$, we can repeat this argument for $[G^{\prime}]_{A}$, and so on, until we obtain a forbidden minor of $G^{\prime}$ for which Statements (i) and (ii) are true.
We complete the proof by showing how to obtain $[G^{\prime}]_{A}$. Let $a^{\prime}$ be the first vertex after $a$ in $P_{a}$ that is not contained in $[a]^{-1}$. Define $b^{\prime}$ similarly. First, we show that we can wlog assume neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. If this is not the case, then assume that $a^{\prime}\in[w]^{-1}$. Let $a^{\prime\prime}$ be the first vertex in $P_{a}$ after $a^{\prime}$ that is not contained in $[w]^{-1}$, and define $S_{a^{\prime}}$ to be the path in $P_{a}$ between $a^{\prime}$ and the vertex that immediately precedes $a^{\prime\prime}$. Let $J$ be the subgraph of $G^{\prime}$ induced by $[w]^{-1}$. Consider the connected component $J_{w}$ of $J\setminus S_{a^{\prime}}$ that contains $w$, and let $J_{a^{\prime}}=J\setminus J_{w}$. Note that $J_{a^{\prime}}$ does not contain any vertex along $P_{b}$ between $w$ and $b$. If $a^{\prime\prime}$ is contained in $[a]^{-1}$ and $J_{w}$ neighbors $[d]^{-1}$, where $[d]$ is the neighbor of $[w]$ other than $[a]$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[w]^{-1}$ to $[a]^{-1}$ that exchanges $J_{a^{\prime}}$ satisfies Statements (i) and (iii). Otherwise, if either $a^{\prime\prime}$ is contained in $[a]^{-1}$ and $J_{w}$ does not neighbor $[d]^{-1}$ or $a^{\prime\prime}$ is contained in $[d]^{-1}$, then the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[w]^{-1}$ to $[d]^{-1}$ that exchanges $J_{a^{\prime}}$ satisfies (i) and $a([G^{\prime}]_{B})=a([G^{\prime}])$ and $b([G^{\prime}]_{B})=b([G^{\prime}])$. By repeating this argument for $[G^{\prime}]_{B}$ and $b^{\prime}$ if necessary, we obtain a minor with the desired property, which we can choose $[G^{\prime}]$ to be to begin with.
Case 1: $[G^{\prime}]$ is $K_{5}$.
If $V_{a}$ is a subset of either $[x]^{-1}$ or $[y]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Else, if $V_{a}$ is a subset of $[d]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (iii). Otherwise, if $V_{a}$ is a subset of $[d]^{-1}\cup[x]^{-1}$ but not a subset $[d]^{-1}$ or $[x]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[y]^{-1}$ that exchanges $H_{b}$ satisfies Statements (i) and (ii). A similar argument works if $V_{a}$ is a subset of $[d]^{-1}\cup[y]^{-1}$ but not a subset $[d]^{-1}$ or $[y]^{-1}$. If $V_{a}$ is a subset of $[x]^{-1}\cup[y]^{-1}$ but not a subset $[x]^{-1}$ or $[y]^{-1}$, then $V_{b}$ contains some vertex in $[d]^{-1}$. Furthermore, using arguments similar to those above, we can assume $V_{b}$ contains some vertex in $[x]^{-1}\cup[y]^{-1}$. Wlog, if this vertex is contained in $[y]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Lastly, the above arguments allow us to assume that $V_{a}$ and $V_{b}$ each contain vertices in $[d]^{-1}$, $[x]^{-1}$, and $[y]^{-1}$, in which case the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii).
Case 2: $[G^{\prime}]$ is $K_{2,2,2}$.
Wlog, assume that the neighborhood of $[a]$ is $\{[d],[e],[w],[x]\}$, where $[e]$ is the vertex in $V([G^{\prime}])\setminus\{[a],[d],[w],[x],[y]\}$. If $V_{a}$ is a subset of $[x]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Else, if $V_{a}$ is a subset of $[d]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (iii). Otherwise, if $V_{a}$ is a subset of $[e]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[e]^{-1}$ that exchanges $H_{a}$. Observe that contracting $[bd]_{B}$ yields a $K_{5}$ minor $[G^{\prime}]_{C}$ in which $[w]_{C}$, $[x]_{C}$, and $[y]_{C}$ are distinct vertices and $wx$ and $wy$ are retained. Hence, the existence of $[G^{\prime}]_{A}$ follows from Case 1. Similar arguments apply for all remaining cases of $V_{a}$.
The above cases are exhaustive, and so the lemma is proved. ∎
Proof.
The proof is similar to the proof of Lemma 56. Let $G^{\prime}=G\cup F$ and assume that $[G^{\prime}]$ exists. Since $G$ is $3$-connected, it contains at least two internally vertex-disjoint paths $P_{a}$ and $P_{b}$ between $w$ and $x$, where $a$ and $b$ are the first vertices along these paths that are not contained in $[w]^{-1}$, respectively. Note that $[x]^{-1}\cup[y]^{-1}\cup[z]^{-1}$ does not contain $a$ or $b$. Also, since $[w]$ has degree one in $[G^{\prime}]\setminus[F]$, we have $[a]=[b]$. We will show how to obtain a minor $[G^{\prime}]_{A}$ from $[G^{\prime}]$ via a sequence of contractions and component exchanges such that $[G^{\prime}]_{A}$ is a forbidden minor in which $[w]_{A}$, $[x]_{A}$, $[y]_{A}$, and $[z]_{A}$ are distinct vertices and either $wx$, $wy$, or $wz$ is doubled. If such a minor exists, then the lemma is clearly proved.
We complete the proof by showing how to obtain $[G^{\prime}]_{A}$. Let $a^{\prime}$ be the first vertex after $a$ in $P_{a}$ that is not contained in $[a]^{-1}$. Define $b^{\prime}$ similarly. First, we show that we can wlog assume neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. If this is not the case, then assume that $a^{\prime}\in[w]^{-1}$. Since $[w]$ has degree one in $[G^{\prime}]\setminus[F]$, the first vertex $a^{\prime\prime}$ after $a^{\prime}$ in $P_{a}$ that is not contained in $[w]^{-1}$ is contained in $[a]^{-1}$. Define $S_{a^{\prime}}$ to be the path in $P_{a}$ between $a^{\prime}$ and the vertex that immediately precedes $a^{\prime\prime}$. Let $J$ be the subgraph of $G^{\prime}$ induced by $[w]^{-1}$. Consider the connected component $J_{w}$ of $J\setminus S_{a^{\prime}}$ that contains $w$, and let $J_{a^{\prime}}=J\setminus J_{w}$. Note that $J_{a^{\prime}}$ does not contain any vertex along $P_{b}$ between $w$ and $b$. The minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[w]^{-1}$ to $[a]^{-1}$ that exchanges $J_{a^{\prime}}$ is a forbidden minor in which $[w]_{B}$, $[x]_{B}$, $[y]_{B}$, and $[z]_{B}$ are distinct vertices and $wx$, $wy$, and $wz$ are retained. By repeating this argument for $[G^{\prime}]_{B}$, and so on, we eventually obtain a minor with the desired property, which we can choose $[G^{\prime}]$ to be to begin with.
Case 1: $[G^{\prime}]$ is $K_{5}$.
If $V_{a}$ is a subset of either $[x]^{-1}$, $[y]^{-1}$, or $[z]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ has the desired properties. Else, if $V_{a}$ is a subset of the union of any two of these sets, wlog say $[x]^{-1}\cup[y]^{-1}$, but not a subset of any one of these sets, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[z]^{-1}$ that exchanges $H_{b}$ has the desired properties. Otherwise, consider the case where $V_{a}$ and $V_{b}$ each contain vertices in $[x]^{-1}$, $[y]^{-1}$, and $[z]^{-1}$, in which case the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ has the desired properties.
Case 2: $[G^{\prime}]$ is $K_{2,2,2}$.
Wlog, assume that the neighborhood of $[a]$ is $\{[d],[w],[x],[y]\}$, where $[d]$ is the vertex in $V([G^{\prime}])\setminus\{[a],[w],[x],[y],[z]\}$. If $V_{a}$ is a subset of either $[x]^{-1}$ or $[y]^{-1}$, then the minor obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ has the desired properties. Else, if $V_{a}$ is a subset of $[d]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$. Contracting $[bx]_{B}$ yields a minor with the desired properties. Otherwise, if $V_{a}$ is a subset of $[x]^{-1}\cup[y]^{-1}$ but not a subset of $[x]^{-1}$ or $[y]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{b}$. Similar arguments apply for all remaining cases of $V_{a}$.
The above cases are exhaustive, and so the lemma is proved. ∎
Proof of Theorem 5.
Assume that $G\cup wu_{i}$ is a partial $3$-tree for each $wu_{i}$. If $G\cup\{wu_{i}\}$ is $3$-flattenaable, then the lemma follows from Lemma 20. Otherwise, let $F$ be a subset of $\{wu_{i}\}$ of minimum size such that there exists a forbidden minor $[G\cup F]$. Consider the case where some pair $f\in F$ is contracted in every forbidden minor of $G\cup F$, and let $F^{\prime}=F\setminus\{f\}$. Proposition 1 shows that $(G\cup F^{\prime},f)$ is a winged graph pair with an $f$-winged graph minor. Therefore, either Proposition 3, if some such minor has singleton wing tips, or Lemma 12, otherwise, shows that $G\cup F^{\prime}$ has a separator of size two. However, this contradicts the fact that $G$ is $3$-connected.
Next, we can assume that some pair in $F$ is preserved in $[G\cup F]$. Let $X$ be the set of all such pairs and let $F^{\prime}=F\setminus X$. The minimality condition on $F$ implies that each pair in $X$ is retained in $[G\cup F]$. Hence, since $G$ is connected and each vertex of any forbidden minor has degree four, we have $|X|\leq 3$. If $|X|>1$, then applying either Lemma 56 or Lemma 57 followed by Lemma 56 to $G\cup F^{\prime}$ and $X$ shows that the minimality condition on $F$ is violated. Therefore, assume that $X$ contains exactly one pair, say $f=[wz]$. Let $H$ be the subgraph of $G$ induced by $[w]^{-1}$ and let $H_{w}$ be its connected component that contains $w$. Also, let $\{[v_{i}]\}$ be the neighborhood of $[w]$ in $[G\cup F]\setminus[z]$. If each set $[v_{i}]^{-1}$ contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$, then $G\cup f$ clearly has a forbidden minor, which contradicts our assumption that $G\cup f$ is $3$-flattenable. Otherwise, if exactly two sets in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$ and $[v_{2}]^{-1}$, each contain some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$, then some connected component $J$ of $H\setminus H_{w}$ contains both some vertex in $\{u_{i}\}$ and some vertex whose neighborhood in $G$ contains some vertex in $[v_{3}]^{-1}$. The minor $[G\cup F]_{A}$ obtained via the component exchange from $[w]^{-1}$ to $[v_{3}]^{-1}$ that exchanges $J$ either is or has a forbidden minor in which at least two pairs in $\{wu_{i}\}$ are preserved, which is a contradiction.
Lastly, assume that exactly one set in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$, contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$. Let $\{J_{i}\}$ be the set of connected components of $H\setminus H_{w}$. Observe that each $J_{i}$ contains some vertex $u_{j}$ and some vertex whose neighborhood in $G$ contains some vertex in at least one of the sets $[v_{i}]^{-1}$ or $[z]^{-1}$. Furthermore, $[v_{2}]^{-1}$ and $[v_{3}]^{-1}$ each contain some vertex whose neighborhood in $G$ contains some vertex in some $J_{i}$. If the set $\{J_{i}\}$ contains more than one component, then wlog let $J_{1}$ contain some vertex whose neighborhood in $G$ contains some vertex in $[v_{2}]^{-1}$. Also, let $x$ be any vertex with $[x]\neq[w]$ and whose neighborhood in $G$ contains some vertex in $J_{2}$. Let $[x]=[v_{3}]$ if possible. We arrive at the same contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to $[x]^{-1}$ that exchanges $J_{2}$. On the other hand, if $J_{1}$ is the only component in $\{J_{i}\}$, then let $y$ be any vertex in $J_{1}$ whose neighborhood in $G$ contains some vertex in $[v_{2}]^{-1}$. Since $G$ is $3$-connected, there exists some vertex $x$ with $[x]\neq[w]$ and whose neighborhood in $G$ contains a vertex in some connected component $K$ of $J_{1}\setminus y$. Let $[x]=[v_{3}]$ if possible. We arrive at the same contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to $[x]^{-1}$ that exchanges $K$. This completes the proof. ∎
Appendix E Results about atoms

### v2

@@ B.3   Proof of Lemmas for Proposition 9
Proof of Lemma 36.
The lemma follows immediately from Lemmas 57-59. ∎
B.4   Proof of Lemma 24 for Proposition 10
Proof of Lemma 24.
We can clearly choose $\ell^{\prime}_{i}$ to satisfy (i) and (ii), and we will show that some such choice satisfies (iii). Let $\{v_{i}\}$ be the vertex set of $G$ with $v_{4}4$ not in $H$ and let $\ell_{1}(v_{1}v_{2})$ and $\ell_{2}(v_{1}v_{2})$ be distinct positive real numbers. Consider the $2$-realizations $p$ and $q$ of $(H,\ell_{1})$ and $(H,\ell_{2})$, respectively, with $p_{i}=p(v_{i})$, $q_{i}=q(v_{i})$, $(0,0)=p_{1}=q_{1}$, $p_{2}=(\ell_{1}(v_{1}v_{2}),0)$, $q_{2}=(\ell_{2}(v_{1}v_{2}),0)$, $p_{3}=(x,y)$ with $y\geq 0$, and $q_{3}=(w,z)$ with $z\geq 0$. Since $\ell_{1}(v_{1}v_{2})$ and $\ell_{2}(v_{1}v_{2})$ are distinct, the line-segment between $p_{2}$ and $q_{2}$ has a perpendicular bisector $L_{1}$. If $p_{3}=q_{3}$, then taking any point $x$ on $L_{1}$ and setting $\ell^{\prime}_{i}(v_{4}v_{j})=\|x-p_{j}\|^{2}$ completes the proof. Otherwise, let $L_{2}$ be the perpendicular bisector of the line-segment between $p_{3}$ and $q_{3}$. If some point $x$ lies on both $L_{1}$ and $L_{2}$, then setting $\ell^{\prime}_{i}(v_{4}v_{j})=\|x-p_{j}\|^{2}$ completes the proof; see Figure 21(a). Else, $L_{1}$ and $L_{2}$ are distinct and parallel; see Figure 21(b). In this case, let $p^{\prime}$ be a realization obtained by rotating $p$ about the origin by an angle of $\theta$, and let $p^{\prime}_{i}=p^{\prime}(v_{i})$. It suffices to show that we can choose $p^{\prime}$ such that the perpendicular bisector of the line-segment between $p^{\prime}_{2}$ and $q_{2}$ intersects the perpendicular bisector of the line-segment between $p^{\prime}_{3}$ and $q_{3}$.
We complete the proof by showing how to choose $\theta$ so that $p^{\prime}$ has the desired properties. Let $p^{\prime}_{i}=p(v_{i})$. It suffices to choose $\theta$ so that the line containing the points $p^{\prime}_{2}$ and $q_{2}$ and the line containing the points $p^{\prime}_{3}$ and $q_{3}$ have distinct slope magnitudes. Note that $p^{\prime}_{2}=(a\cos\theta,-a\sin\theta)$, where $a=\ell_{1}(v_{1}v_{2})$, and $p^{\prime}_{3}=(x\cos\theta+y\sin\theta,-x\sin\theta+y\cos\theta)$. Let $a^{\prime}=\ell_{2}(v_{1}v_{2})$. Then, $\theta$ must not be a solution to
$\left|\frac{-a\sin\theta}{a\cos\theta-a^{\prime}}\right|=\left|\frac{-x\sin\theta+y\cos\theta-z}{x\cos\theta+y\sin\theta-w}\right|.$
This is equivalent to $\theta$ not being in the union of the solution sets to any of the following:
$\left|a\cos\theta-a^{\prime}\right|=0,$
$\left|x\cos\theta+y\sin\theta-w\right|=0,$
$\left|(az+a^{\prime}y)\cos\theta+(aw-a^{\prime}x)\sin\theta-(ay+a^{\prime}z)\right|=0,$
$\left|(az+a^{\prime}y)\cos\theta-(aw+a^{\prime}x)\sin\theta+2ax\cos\theta\sin\theta-ay\cos 2\theta-a^{\prime}z\right|=0.$
Each of these equations is periodic and has finitely many solutions on the interval $[0,2\pi]$, and thus we can choose $\theta$ as needed. ∎
Appendix C Results about atoms